# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ 4.1. $L^{2}$-uniform bound of the logarithmic derivative $\mathsf{d}^{N}$ of $\mu^{N}$
4.2. $L^{2}$-uniform bound of logarithmic derivative of $\mu_{R}^{N}$.
Let $\mu_{R}^{N}=\mu^{N}\circ\pi_{\overline{S}_{R}}^{-1}$, where $\pi_{\overline{S}_{R}}(\cdot)=\mathsf{s}(\cdot\cap\overline{S}_{R})$. Let $\overline{\mathsf{S}}_{R}$ is the configuration space over $\overline{S}_{R}$. Then, $\mu_{R}^{N}$ can be regarded as a probability measure on $\overline{\mathsf{S}}_{R}$. Let $\mu_{R}^{N,[1]}$ be the one-reduced Campbell measure of $\mu_{R}^{N}$ defined by $d\mu_{R}^{N,[1]}=\rho^{N,1}({x})d\mu_{R,{x}}^{N}d{x}$. Here, $\mu_{R,{x}}^{N}$ is the reduced Palm measure of $\mu_{R}^{N}$ conditioned at x{x} and $\rho^{N,1}$ is the one-point correlation function of $\mu_{R}^{N}$. Note that the one-point correlation function of $\mu_{R}^{N}$ coincides with that of $\mu^{N}$ on $\overline{S}_{R}$. We can regard $\mu_{R}^{N,[1]}$ as a measure on $\overline{S}_{R}\times\overline{\mathsf{S}}_{R}$.
Let $\mathsf{d}_{R}^{N}$ is the logarithmic derivative of $\mu_{R}^{N}$, that is,
$\int_{\overline{S}_{R}\times\overline{\mathsf{S}}_{R}}\mathsf{h}\mathsf{d}_{R}^{N}d\mu_{R}^{N,[1]}=-\int_{\overline{S}_{R}\times\overline{\mathsf{S}}_{R}}\nabla_{{x}}\mathsf{h}d\mu_{R}^{N,[1]}$ (4.10)
for each $\mathscr{F}_{R}^{[1]}$-measurable $\mathsf{h}\in C_{0}^{\infty}(S_{R})\otimes\mathscr{D}_{\circ\mathrm{b}}$, where $\mathscr{F}_{R}^{[1]}=\mathscr{B}(\overline{S}_{R})\times\mathscr{F}_{R}$ and $\mathscr{F}_{R}=\sigma[\pi_{\overline{S}_{R}}]$. Thus, we have generalized the concept of the logarithmic derivative to the random point field $\mu_{R}^{N}$ on $\overline{S}_{R}$.
We regard $\mathsf{d}_{R}^{N}$ as an $\mathscr{F}_{R}^{[1]}$-measurable function on $\mathbb{R}^{d}\times\mathsf{S}$ such that $\mathsf{d}_{R}^{N}({x},\mathsf{s})=0$ for ${x}\notin\overline{S}_{R}$ and consider $\mu_{R}^{N,[1]}$ to be a measure on $(\mathbb{R}^{d}\times\mathsf{S},\mathscr{B}(\mathbb{R}^{d})\times\mathscr{F}_{R})$ such that $\mu_{R}^{N,[1]}(S_{R}^{c}\times\mathsf{S})=0$. Then, (4.10) can be written as
$\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\mathsf{d}_{R}^{N}d\mu^{N,[1]}=-\int_{\mathbb{R}^{d}\times\mathsf{S}}\nabla_{{x}}\mathsf{h}d\mu^{N,[1]}$ (4.11)
for each $\mathscr{F}_{R}^{[1]}$-measurable $\mathsf{h}\in C_{0}^{\infty}(S_{R})\otimes\mathscr{D}_{\circ\mathrm{b}}$. We often use the second interpretation because it allows us to regard $\mathsf{d}_{R}^{N}({x},\mathsf{s})$ as a function defined on the same set $\mathbb{R}^{d}\times\mathsf{S}$ independent of $R\in\mathbb{N}$.
Remark 4.1.
If we replace $C_{0}^{\infty}(S_{R})\otimes\mathscr{D}_{\circ\mathrm{b}}$ by $C^{\infty}(\overline{S}_{R})\otimes\mathscr{D}_{\circ\mathrm{b}}$ as the space of test functions in (4.10), then a local-time-type singular drift appears in the representation of $\mathsf{d}_{R}^{N}$, indicating the reflection of particles on the boundary $\partial S_{R}$.
Let $\mathfrak{m}^{N}$ be the labeled density of $\mu^{N}$ with respect to the Lebesgue measure on $(\mathbb{R}^{d})^{N}$, where $\mathfrak{m}^{N}$ is given by (2.18). Then, $\mu_{R}^{N,[1]}$ has the labeled density $\mathfrak{m}_{R}^{N,[1]}$ defined on $\overline{S}_{R}\times\sum_{k=0}^{N-1}\overline{S}_{R}^{k}$ such that, for $0\leq k\leq N-1$ and k+l=N1k+l=N-1,
$\mathfrak{m}_{R}^{N,[1]}({x},\mathbf{s})$ $=\frac{1}{\mathscr{Z}_{R}^{N,k}({x})}\int_{(\overline{S}_{R}^{c})^{l}}\mathfrak{m}^{N}({x},\mathbf{s},\mathbf{t})d\mathbf{t},\quad\text{ for }\mathbf{s}\in\overline{S}_{R}^{k},$
$\mathscr{Z}_{R}^{N,k}({x})$ $=\int_{\overline{S}_{R}^{k}\times(\overline{S}_{R}^{c})^{l}}\mathfrak{m}^{N}({x},\mathbf{s},\mathbf{t})d\mathbf{s}d\mathbf{t}\quad\text{ for }{x}\in\overline{S}_{R}.$
Because $\mathfrak{m}^{N}(\mathbf{{x}})$ is a symmetric function of $\mathbf{{x}}=({x}^{i})_{i=1}^{N}$, $\mathfrak{m}_{R}^{N,[1]}$ is well defined.
The logarithmic derivate $\mathsf{d}_{R}^{N}\!:\!\overline{S}_{R}\times\overline{\mathsf{S}}_{R}\!\to\!\mathbb{R}^{d}$ is given by $\mathsf{d}_{R}^{N}({x},\mathsf{s})=0$ for $\mathsf{s}(\overline{S}_{R})\geq N$ and it does by
@@
$\mathsf{d}_{R}^{N}({x},\mathsf{s})=\nabla_{{x}}\log\mathfrak{m}_{R}^{N,[1]}({x},\mathbf{s})$ (4.12)
for $\mathsf{s}(\overline{S}_{R})\leq N-1$, where $\mathsf{s}=\sum_{i=1}^{k}\delta_{s^{i}}$ is defined by $\mathbf{s}=(s^{i})_{i=1}^{k}\in\overline{S}_{R}^{k}$.
Let $0\leq k\leq N-1$ and k+l=N1k+l=N-1. Let $\mathbf{t}\in(\overline{S}_{R}^{c})^{l}$. We set, for $\mathbf{s}=(s^{i})_{i=1}^{k}\in\overline{S}_{R}^{k}$ and $\mathsf{s}=\sum_{i=1}^{k}\delta_{s^{i}}\in\overline{\mathsf{S}}_{R}$,
where $f\!:\!\overline{S}_{R}\times\mathsf{S}^{N-1}\!\to\!\mathbb{R}$, $\mathsf{S}^{N-1}=\{\mathsf{u}\in\mathsf{S};\mathsf{u}(\mathbb{R}^{d})=N-1\}$, and $\check{f}\!:\!\overline{S}_{R}\times(\mathbb{R}^{d})^{N-1}\!\to\!\mathbb{R}$ is the function such that $\check{f}({x},\mathbf{s},\mathbf{t})=f({x},\mathsf{s})$ and that $\check{f}({x},\mathbf{s},\mathbf{t})$ is symmetric in $(\mathbf{s},\mathbf{t})$.
The following identity plays an important role in our analysis.
Lemma $4.$
Let $\mathsf{d}^{N}$ and $\mathsf{d}_{R}^{N}$ be as in (4.1) and (4.12), respectively. Then,
@@
Let k+l=N1k+l=N-1. Let $\mathsf{s}$, $\mathbf{s}$, and $\mathbf{t}$ be as in (4.13). Then,
@@
Thus, we obtain (4.14). ∎
As we explained around (4.11), we consider $\mathsf{d}_{R}^{N}$ as a function on $\mathbb{R}^{d}\times\mathsf{S}$. We now obtain a uniform $L^{2}$-bound of the logarithmic derivatives $\mathsf{d}_{R}^{N}$ of $\mu_{R}^{N}$.
@@
Assume ((A1))((A3)). Let $\mathsf{f}_{Q}$ be as in (4.5). Let $R\geq Q+1$. Then,
@@
Recall that $\mu^{N,[1]}(d{x}d\mathsf{s})=\rho^{N,1}({x})\mu^{N}_{x}(d\mathsf{s})d{x}$ and $0\leq\mathsf{f}_{Q}({x})\leq 1$. Then,
$\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{f}_{Q}|\mathsf{d}_{R}^{N}|^{2}d\mu^{N,[1]}$ $=\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{f}_{Q}\big{|}\langle\mathsf{d}^{N}\rangle_{R}^{N}\big{|}^{2}d\mu^{N,[1]}$ by Lemma 4.4
$\leq\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{f}_{Q}\langle|\mathsf{d}^{N}|^{2}\rangle_{R}^{N}d\mu^{N,[1]}$ by Jensen’s inequality
@@
Hence, we obtain (4.15) from Lemma 4.3. ∎
4.3. Strong convergence of density functions of $\mu_{R}^{N}$.
Let $\Lambda_{R}$ be the Poisson random point field with intensity measure $1_{S_{R}}d{x}$. Let $\mu_{R}=\mu\circ\pi_{\overline{S}_{R}}^{-1}$. Let $\Lambda_{R}^{[1]}$ and $\mu_{R}^{N,[1]}$ be the one-reduced Campbell measures of $\Lambda_{R}$ and $\mu_{R}^{N}$, respectively. We denote by $\mathfrak{m}_{R}^{N,[1]}$ the Radon–Nikodym density of $\mu_{R}^{N,[1]}$ with respect to $\Lambda_{R}^{[1]}$. The existence of $\mathfrak{m}_{R}^{N,[1]}$ is obvious from (2.18). Let
@@
Let $\overline{\mathsf{S}}_{R}$ be the configuration space over $\overline{S}_{R}$. Then, we naturally regard $\overline{\mathsf{S}}_{R,k}^{[1]}$ as a subset as a subset of $\overline{S}_{R}\times\overline{\mathsf{S}}_{R}$ as follows: $\overline{\mathsf{S}}_{R,k}^{[1]}=\{({x},\mathsf{s})\in\overline{S}_{R}\times\overline{\mathsf{S}}_{R};\,\mathsf{s}(\overline{S}_{R})=k\}$.
@@
Let $\mathsf{f}_{Q}$ be as in (4.5). Let $Q,R\in\mathbb{N}$ such that $R\geq Q+1$. Then, for all $k\in\{0\}\cup\mathbb{N}$ and $l\in\mathbb{N}$, the following hold.
i $\{\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{N_{\mathfrak{n}},[1]}\}_{\mathfrak{n}\in\mathbb{N}}$ is relatively compact in $L^{2}(\Lambda_{R}^{[1]})$.
ii There exists a family of non-negative functions $\mathfrak{m}_{R,k,l}^{[1]}$ such that
$\lim_{\mathfrak{n}\to\infty}\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{N_{\mathfrak{n}},[1]}=\mathsf{f}_{Q}\mathfrak{m}_{R,k,l}^{[1]}\text{ in }L^{2}(\Lambda_{R}^{[1]}),$ (4.17)
$\mathfrak{m}_{R,k,l}^{[1]}=\mathfrak{m}_{R,k,l+1}^{[1]}\wedge l.$ (4.18)

### v4

@@ 5.2. Explicit formulae of the logarithmic derivative $\mathsf{d}_{R,\mathsf{y}}$ of $\mu_{R,\mathsf{y}}$
[…]
The convergence above holds in $C^{1}(\overline{S}_{R,\epsilon}(\mathsf{s}))$ for each $\epsilon>0$ and all $\mathsf{s}$ such that $\mathsf{s}=\sum_{s^{i}\in\overline{S}_{R}}\delta_{s^{i}}$, where $\overline{S}_{R,\epsilon}(\mathsf{s})=\{{x}\in\overline{S}_{R};\lvert{x}-s^{i}\rvert\geq\epsilon\text{ for all }i\}$.
By Lemma 5.5 and the above equation, we obtain, weakly in $L^{1}(\Lambda_{R}^{[1]})$,
$\lim_{\mathfrak{n}\to\infty}\Big(\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{y}}^{\mathbf{i},\mathfrak{n},\mathfrak{n}}{x}^{\mathbf{i}}\Big)\mathsf{f}_{Q}\mathfrak{m}_{R,\mathsf{y}}^{[1]}=\lim_{\mathfrak{n}\to\infty}\Big(\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{y}}^{\mathbf{i},\mathfrak{n},\mathfrak{n}}{x}^{\mathbf{i}}\Big)\mathsf{f}_{Q}\mathfrak{m}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[1]}$
$=\lim_{\mathfrak{n}\to\infty}\Big(\Big(\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{y}}^{\mathbf{i},\mathfrak{n},\mathfrak{n}}{x}^{\mathbf{i}}\Big)-\mathsf{d}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}}}({x},\mathsf{s})+\mathsf{d}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}}}({x},\mathsf{s})\Big)\mathsf{f}_{Q}\mathfrak{m}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[1]}$
$=\Big(\Big(-\mathscr{R}_{R,\mathsf{y}}^{\ell}({x})+\nabla\Phi({x})+\sum_{s^{i}\in\overline{S}_{R}}\nabla_{{x}}\Psi({x}-s^{i})\Big)+\mathsf{d}_{R,\mathsf{y}}({x},\mathsf{s})\Big)\mathsf{f}_{Q}\mathfrak{m}_{R,\mathsf{y}}^{[1]}.$
From $\mu_{R,\mathsf{y}}^{[1]}(\overline{S}_{R}\times\overline{\mathsf{S}}_{R})>0$, we have $\int_{\overline{S}_{R}\times\overline{\mathsf{S}}_{R}}\mathfrak{m}_{R,\mathsf{y}}^{[1]}d\Lambda_{R}^{[1]}>0$. Hence, the family $\{{x}^{\mathbf{i}}\mathsf{f}_{Q}({x},\mathsf{s})\mathfrak{m}_{R,\mathsf{y}}^{[1]}({x},\mathsf{s});\mathbf{i}\in\mathbb{I}(\ell)\}$ is linearly independent in $L^{1}(\Lambda_{R}^{[1]})$.
From this and the display equation right above, each numerical sequence $\{\eta_{R,\mathsf{y}}^{\mathbf{i},\mathfrak{n},\mathfrak{n}}\}_{\mathfrak{n}\in\mathbb{N}}$ has a finite limit $\eta_{R,\mathsf{y}}^{\mathbf{i}}$ satisfying
$\lim_{\mathfrak{n}\to\infty}\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{y}}^{\mathbf{i},\mathfrak{n},\mathfrak{n}}\big({x}^{\mathbf{i}}\mathsf{f}_{Q}\mathfrak{m}_{R,\mathsf{y}}^{[1]}\big)=\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{y}}^{\mathbf{i}}\big({x}^{\mathbf{i}}\mathsf{f}_{Q}\mathfrak{m}_{R,\mathsf{y}}^{[1]}\big),$
which implies the first equality in (5.25).
Let ${x}_{\mathbf{j}}^{\mathbf{i}}$ for ${x}\in\mathscr{S}_{R}^{\ell}$ be as in (2.15). For each $\mathbf{j}\in\mathbb{I}(\ell)$,
[…]
Let $\mathbf{x}_{\mathbf{j}}=({x}_{\mathbf{j}}^{\mathbf{i}})_{\mathbf{i}\in\mathbb{I}(\ell)}$ for $\mathbf{j}\in\mathbb{I}(\ell)$. Then the vectors $\{\mathbf{x}_{\mathbf{j}}\}_{\mathbf{j}\in\mathbb{I}(\ell)}$ are linearly independent in $\mathbb{R}^{\mathbb{I}(\ell)}$ for each ${x}\in\mathscr{S}_{R}^{\ell}$ by (2.15). Hence, using the display equation right above, we find that, for each $\mathbf{i}\in\mathbb{I}(\ell)$, the numerical sequence $\eta_{R,\mathsf{y}}^{\mathbf{i},\mathfrak{n}}-\eta_{R,\mathsf{y}}^{\mathbf{i},\mathfrak{n},\mathfrak{n}}$ converges to zero as $\mathfrak{n}\to\infty$ for $\mu$-a.s. $\mathsf{y}$. From this and $\lim_{\mathfrak{n}\to\infty}\eta_{R,\mathsf{y}}^{\mathbf{i},\mathfrak{n},\mathfrak{n}}=\eta_{R,\mathsf{y}}^{\mathbf{i}}<\infty$ for each $\mathbf{i}\in\mathbb{I}(\ell)$, we obtain (5.25).
Clearly, $\eta_{R,\mathsf{y}}^{\mathbf{i},\mathfrak{n}}$ and $\eta_{R,\mathsf{y}}^{\mathbf{i},\mathfrak{n},\mathfrak{n}}$ are independent of $\pi_{\overline{S}_{R}}(\mathsf{y})$. Hence, (5.25) yields (5.26). From (5.22) and (5.25), we have
@@
$\eta_{R,\mathsf{y}}^{\mathbf{i}}=\lim_{\mathfrak{n}\to\infty}\eta_{R,\mathsf{y}}^{\mathbf{i},\mathfrak{n}}=\Big(\frac{1}{i_{1}!\cdots i_{d}!}\Big)\lim_{\mathfrak{n}\to\infty}\sum_{{y}^{i}\not\in\overline{S}_{R}}^{N_{\mathfrak{n}}}(-\nabla\Psi)^{(\mathbf{i})}(0-{y}^{i}).$
Because the sum in the display equation right above converges, we deduce
This implies (5.27). From (5.23), (5.25), and (5.18), we have (5.28). ∎
Let $\overline{\mathsf{S}}_{R}$ be the configuration space over $\overline{S}_{R}$. Let
$\overline{\mathsf{S}}_{R,\epsilon}^{[1],k}=\{({x},\mathsf{s})\in\overline{S}_{R}\times\overline{\mathsf{S}}_{R}\,;\,\lvert{x}-s^{i}\rvert\geq\epsilon,\ \mathsf{s}(\overline{S}_{R})=k\}.$ (5.29)
We regard $f({x},\mathsf{s})\in C^{1}(\overline{\mathsf{S}}_{R,\epsilon}^{[1],k})$ as a function of the variables $({x},s^{i})_{i=1}^{k}$, symmetric in $(s^{i})_{i=1}^{k}\in\overline{S}_{R}^{k}$, and equip $C^{1}(\overline{\mathsf{S}}_{R,\epsilon}^{[1],k})$ with the $C^{1}$-norm. Here $C^{1}(\overline{\mathsf{S}}_{R,\epsilon}^{[1],k})=C^{1}(S_{R})$ for $k=0$ by convention.
Theorem 5.1.
Let $\mathsf{d}_{R,\mathsf{y}}$ be defined in Lemma 5.5.
i For each $R\in\mathbb{N}$ and for $\mu$-a.s. $\mathsf{y}$, the identity
$\mathsf{d}_{R,\mathsf{y}}=\beta\Big\{-\nabla\Phi({x})-\sum_{s^{i}\in\overline{S}_{R}}\nabla\Psi({x}-s^{i})+\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{y}}^{\mathbf{i}}{x}^{\mathbf{i}}+\mathscr{R}_{R,\mathsf{y}}^{\ell}({x})\Big\}$ (5.30)
holds for all $({x},\mathsf{s})\in S_{R}\times\mathsf{S}$ such that ${x}\neq s^{i}$ for all $i$.
ii For $\mu$-a.s. $\mathsf{y}$ and each $R,k\in\mathbb{N}$ and each $\epsilon>0$, we have
$\mathsf{d}_{R,\mathsf{y}}({x},\mathsf{s})=\lim_{\mathfrak{n}\to\infty}\mathsf{d}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}}}({x},\mathsf{s})\quad\text{in }C^{1}(\overline{\mathsf{S}}_{R,\epsilon}^{[1],k}).$ (5.31)
Here $N_{\mathfrak{n}}$ and $\mathsf{d}_{R,\mathsf{y}^{N}}^{N}$ are as in ((A2))iii and (5.6), respectively.
@@
Let $\mathsf{S}_{R}^{m}=\{\mathsf{s}\in\mathsf{S}\,;\,\mathsf{s}(S_{R})=m\}$. Let $\mathsf{h}=\mathsf{f}\otimes\mathsf{g}\in C_{0}^{\infty}(S_{R})\otimes\mathscr{D}_{\circ\mathrm{b}}$ be an $\mathscr{F}_{R}^{[1]}$-measurable function such that $\mathsf{g}(\mathsf{s})=0$ on $\bigsqcup_{m=M}^{\infty}\mathsf{S}_{R}^{m}$ for some $M\in\mathbb{N}$, and $\mathsf{h}({x},\mathsf{s})=0$ whenever $\lvert{x}-s^{i}\rvert\leq\epsilon$ for some ii and some $\epsilon>0$. Then
$\Big(\sum_{s^{i}\in\overline{S}_{R}}\nabla\Psi({x}-s^{i})\Big)\mathsf{h}({x},\mathsf{s})\in C_{0}^{\infty}(S_{R})\otimes\mathscr{D}_{\circ\mathrm{b}}.$ (5.32)
Let $\mu_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[1]}$ and $\mathsf{y}^{N_{\mathfrak{n}}}$ be as in (5.1) and (2.23) with $N=N_{\mathfrak{n}}$, respectively. By (2.4), (2.7), (5.9), and (5.32), we obtain
Hence, we obtain
$=$ $\lim_{\mathfrak{n}\to\infty}\int_{\mathbb{R}^{d}\times\mathsf{S}}\Big(\mathsf{d}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}}}({x},\mathsf{s})+$
@@
[…]
This yields (5.30) for $\mu_{R,\mathsf{y}}^{[1]}$-a.s. $({x},\mathsf{s})$. The right-hand side of (5.30) is continuous on $\{({x},\mathsf{s})\in\overline{S}_{R}\times\overline{\mathsf{S}}_{R};{x}\neq s^{i}\}$. Hence, there exists a $\mu_{R,\mathsf{y}}^{[1]}$-version of $\mathsf{d}_{R,\mathsf{y}}$ such that (5.30) holds for all $\{({x},\mathsf{s})\in\overline{S}_{R}\times\overline{\mathsf{S}}_{R};{x}\neq s^{i}\}$. This yields i.
For $\mu$-a.s. $\mathsf{y}$, by (5.6) and (5.23), we have, ${x}\in O_{N_{\mathfrak{n}}}$,
Hence, by (2.4) and (2.7), together with Propositions 5.1 and 5.2, we obtain
$\lim_{\mathfrak{n}\to\infty}\mathsf{d}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}}}({x},\mathsf{s})$ $=\beta\Big\{-\nabla\Phi({x})-\sum_{s^{i}\in\overline{S}_{R}}\nabla\Psi({x}-s^{i})+\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{y}}^{\mathbf{i}}{x}^{\mathbf{i}}+\mathscr{R}_{R,\mathsf{y}}^{\ell}({x})\Big\}.$
in $C^{1}(\overline{\mathsf{S}}_{R,\epsilon}^{[1],k})$ for each $\epsilon>0$. From this and i, we obtain ii. ∎
5.3. Explicit formulas for $\mathsf{d}^{\mu_{\mathsf{y}}}$ and $\mathsf{d}^{\mu}$: Proof of Theorem 2.1
We continue to work under Assumptions ((A1))((A3)).
Let $\mathsf{d}_{R,\mathsf{y}}$ be as in Lemma 5.5. The consistency of $\mathsf{d}_{R,\mathsf{y}}$ in the following lemma is crucial for demonstrating the explicit expression of the logarithmic derivatives of $\mu$ in Theorem 2.1.
@@
For $\mu$-a.s. $\mathsf{y}$ and each $R\in\mathbb{N}$, we have the following:
@@
for all $({x},\mathsf{s})\in\overline{S}_{R}\times\overline{\mathsf{S}}_{R}$ such that ${x}\neq s^{i}$ for each $i$.
@@
From (5.6), we have
@@
Hence, applying (5.31) to (5.35), we obtain (5.34). ∎
Let $\mathsf{S}_{0}$ be as in Definition 2.2. We introduce the equivalence relation in $\mathsf{S}_{0}$ such that $\mathsf{a}\sim_{\mathscr{T}}\mathsf{b}$ if and only if (2.37) holds.
Let $\mathsf{y}]=\{\mathsf{z}\in\mathsf{S}_{0};\mathsf{y}\sim_{\mathscr{T}}\mathsf{z}\}$. Let $\mathsf{S}^{[1]}[\mathsf{y}]=\{({x},\mathsf{s});\delta_{x}+\mathsf{s}\in[\mathsf{y}]\}$ and
$\mathsf{S}_{\neq}^{[1]}[\mathsf{y}]=\{({x},\mathsf{s})\in\mathsf{S}^{[1]}[\mathsf{y}];x\neq s^{i}\text{ for all }i,\ \mathsf{s}\in\mathsf{S}_{\mathrm{s}}\},\quad\mathsf{s}=\sum_{i}\delta_{s^{i}},$ (5.36)
where $\mathsf{S}_{\mathrm{s}}=\{\mathsf{s}\in\mathsf{S}\,;\,\mathsf{s}(\{{x}\})\in\{0,1\}\text{ for all }{x}\in\mathbb{R}^{d}\}$.
For $\mu$-a.s. $\mathsf{y}$, we define the function $\mathfrak{d}_{\mathsf{y}}({x},\mathsf{s})$ on $\mathsf{S}_{\neq}^{[1]}[\mathsf{y}]$ by
@@
$\mathfrak{d}_{\mathsf{y}}({x},\mathsf{s})=\mathfrak{d}_{\mathsf{z}}({x},\mathsf{s})\quad\text{ for all }({x},\mathsf{s})\in\mathsf{S}_{\neq}^{[1]}[\mathsf{y}],\,\mathsf{z}\in[\mathsf{y}].$
We also set $\mathfrak{d}_{\mathsf{y}}({x},\mathsf{s})=0$ for $\mathsf{y}\notin\mathsf{S}_{0}$. Because the equivalence relation $\sim_{\mathscr{T}}$ gives the partition of $\mathsf{S}_{0}$ and $\mu(\mathsf{S}_{0})=1$, such a version of $\mathfrak{d}_{\mathsf{y}}$ is well defined for all $({x},\mathsf{s})\in\mathbb{R}^{d}\times\mathsf{S}$. Thus, for any $({x},\mathsf{s})\in\mathbb{R}^{d}\times\mathsf{S}$, we define $\mathfrak{d}$ by
$\mathfrak{d}({x},\mathsf{s})=\mathfrak{d}_{\mathsf{y}}({x},\mathsf{s})\quad\text{for }({x},\mathsf{s})\in\mathsf{S}_{\neq}^{[1]}[\mathsf{y}],\,\mathsf{y}\in\mathsf{S}.$ (5.38)
Proposition 5.3.
i For $\mu$-a.s. $\mathsf{y}$, $\mathfrak{d}$ is the $\bullet$-logarithmic derivative $\mathsf{d}^{\mu_{\mathsf{y}}}$ of $\mu_{\mathsf{y}}$.
$i$ $\mathfrak{d}$ is the $\bullet$-logarithmic derivative $\mathsf{d}^{\mu}$ of $\mu$ and satisfies (2.26).
Proof.
From (5.12) and (5.37), we obtain that, for all $\mathsf{h}\in\mathscr{D}_{\circ\mathrm{b}}^{[1]}$,
$\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\,\mathfrak{d}_{\mathsf{y}}\,d\mu_{\mathsf{y}}^{[1]}=-\int_{\mathbb{R}^{d}\times\mathsf{S}}\nabla_{{x}}h\,d\mu_{\mathsf{y}}^{[1]}.$ (5.39)
By Lemma 13.6, (5.39) holds for all $\mathsf{h}\in\mathscr{D}_{\bullet\mathrm{b}}^{[1]}$, and hence $\mathfrak{d}_{\mathsf{y}}$ is the $\bullet$-logarithmic derivative of $\mu_{\mathsf{y}}$. Combining this with (5.38), we obtain i.
Using $\mu^{[1]}=\int_{\mathsf{S}}\mu_{\mathsf{y}}^{[1]}\,\mu(d\mathsf{y})$ and (5.39), we further obtain, for all $\mathsf{h}\in\mathscr{D}_{\bullet\mathrm{b}}^{[1]}$,
$\int_{\mathbb{R}^{d}\times\mathsf{S}}\mathsf{h}\,\mathfrak{d}\,d\mu^{[1]}=-\int_{\mathbb{R}^{d}\times\mathsf{S}}\nabla_{{x}}h\,d\mu^{[1]}.$
Hence, $\mathfrak{d}$ is the $\bullet$-logarithmic derivative of $\mu$.
Finally, from this, (5.37), and (5.38), for $\mu$-a.s. $\mathsf{y}$ and $\mu^{[1]}_{\mathsf{y}}-$a.e. $({x},\mathsf{s})$,
$\mathsf{d}^{\mu}({x},\mathsf{s})=\mathsf{d}_{R,\mathsf{s}}\bigl({x},\pi_{\overline{S}_{R}}(\mathsf{s})\bigr).$
Together with (5.30), this yields (2.26) and completes the proof of ii. ∎
Proof of Theorem 2.1. We obtain the $\bullet$-logarithmic derivative $\mathsf{d}^{\mu}\in L_{\mathrm{loc}}^{2}(\mu^{[1]})$ denoted by (2.26) from Propositions 4.2 and 5.3. We will prove the remaining claims of i later.
From (5.27), $\lim_{R\to\infty}\eta_{R,\mathsf{s}}^{\mathbf{i}}=0$. From (5.17), $\lim_{R\to\infty}\mathscr{R}_{R,\mathsf{s}}^{\ell}=0$ in $C^{1}(\overline{S}_{Q})$. Hence, we deduce (2.27). From (2.26) and (2.27), we have, for $\mu$-a.s. $\mathsf{s}$,
$\mathsf{d}^{\mu}$ $({x},\mathsf{s})=\beta\Big(-\nabla\Phi({x})-\sum_{s^{i}\in\overline{S}_{R}}\nabla\Psi({x}-s^{i})+\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{s}}^{\mathbf{i}}{x}^{\mathbf{i}}+\mathscr{R}_{R,\mathsf{s}}^{\ell}({x})\Big)$
[…]
for all $Q,R\in\mathbb{N}$ and each $\epsilon>0$. Hence, we obtain (2.28). Thus, we have obtained ii.
In addition to ((A1))((A3)), we assume ((A4)) in the proof of iii. From (2.28) and (2.20), it follows that, for $\mu^{[1]}$-a.e. $x$,
[…]
$-\beta\lim_{R\to\infty}\sum_{i}1_{S_{R}}({x}-s^{i})\nabla\Psi({x}-s^{i})$
[…]
This proves iii. ∎
As mentioned before ((A3)), ((A3)) is automatically satisfied when $\Psi^{N}=\Psi$. The following lemma verifies the only nontrivial condition in ((A3)).
@@
Let $\ell_{0}$ be as in ((A2)). Let $d+\ell>\ell_{0}$. Then (2.17) holds.